The dielectric strength of air (that is, the maxim
Answer #1

given electric strength is 3 *10^6

radius of the spherical conductor is 30 cms?

we have to find th max charge placed on the sphere before dielectric breakdown occur?

for this charge what is the electric potential?

E = 3.0×10^6 V/m
The permittivity of air is approximately:
εo ≈ 8.85*10^-12 C^2.N-1.m-2
r = 0.3m

Via Gauss's law, the the maximum charge is:
∫s E.dA = Q/εo
= E*4πr^2 = Q/εo
Q = E*4πr^2*εo
Q = 3*10^6 *4*3.14*0.3*0.3*8.85*10^-12

Q= 30.01*10-^6 on sphere surface:

chargeQ = 30 micro coumbs

now potential

V = Q/(4πεo*r)

V = 30.01*10-^6 /4*3.14*8.85*10^-12*0.3

V = 0.9 *10^6N/C